solve the equation
10x^2-2x=0

Answer by RAY100(1637)

10x^2-2x=0
2x(5x-1)=0
2x=0,,x=0
5x-1=0,,,5x=1,,,x=1/5
checking
for x=0, 0-0=0,,,ok
for x=1/5,,,,10*(1/5)^2-2(1/5)=0
10/25 -2/5=0
10/25-10/25=0 ok